# Belinda Z. Li    Zifan Carl Guo    Jacob Andreas

### v2

@@ 4 What Mechanisms do Transformers Learn?
Figure 2: Activation patching on the residual stream for various Pythia models trained on $S_{3}$ and $S_{5}$. Each cell at layer $l$ and token $t$ represents the probability of the correct final state when the entire prefix up to $t$ at layer $l$ is restored. We find signatures matching the AA and PAA algorithms from Figure 1, with both models ignoring exponentially longer prefixes as we traverse down the layers, and PAA models having a middle chunk containing some information about the final state, but not its parity.
Nevertheless, our experiments successfully rule out some possible state tracking mechanisms, and identify algorithmic features likely to be shared between the idealized mechanisms above and the true behavior learned by transformers. Specifically, our experiments yield evidence consistent with the associative algorithm (AA) in some models and the parity-associative algorithm (PAA) in other models, across architectures, sizes, and initializations.
@@ 4.1 Experimental Setup
We generate 1 million unique length-100 sequences of permutations in both $S_{3}$ and $S_{5}$. We split the data 90/10 for training/analysis. fine-tune these models (using a cross-entropy loss) to predict the state corresponding to each prefix of each action sequence:
@@
where $p_{\texttt{LM}}\left(s_{n}\mid a_{0}\ldots a_{t}\right)$ is the probability the language model places on state token $s_{n}$ when conditioned on the lenght-$n$ prefix of the document.
@@
Figure 3: Accuracy of state probe and state parity probe across layers on $S_{3}$ and $S_{5}$ models sometimes match signatures for AA, and sometimes PAA. In all models, the state probe accuracy increases roughly exponentially with model depth. We find that in PAA models, the parity of the state is linearly decodeable from earlier intermediate layers, while in the AA models shown above, the parity is never linearly encoded in any layer of the model. (In other AA models, the parity can only be linearly decoded at the final layer.)
Figure 4: In models that learn PAA on $S_{3}$, representations of the final product can be geometrically decomposed into two orthogonal directions, corresponding to the parity of the product (represented as the Z-axis in the above graph) and cluster identity of the product (represented by the X-Y plane). Note that the clusters are at 60 degrees at each other and products of different parities within a cluster are equidistant from each other, with odd-parity products in one plane, and even-parity products in another plane.
@@ 4.2 Activation Patching
For both the $S_{3}$ and $S_{5}$ tasks, across training runs, we find that activation patching results exhibit two broad clusters of behavior. For some trained models, they match the activation patching signature associated with AA; in others, they match the signature of PAA—even when the only source of variability across training runs is the order in which data is presented. Results for prototypical AA- and PAA-type models, on both $S_{3}$ and $S_{5}$, are shown in Figure 2. Additional patching results can be found in Appendix B, confirming that patching intermediate representations of PAA-type models (the light-colored cells in Figure 2) results specifically in predictions with incorrect parity, as predicted.
@@ 4.3 Probing
Test set accuracies of linear probes across LM layers $l$ are plotted in Figure 3. We again find empirical signatures consistent with those predicted by AA and PAA, on both $S_{3}$ and $S_{5}$. Models with AA-type probing signatures always have AA-type patching signatures, and vice-versa. Throughout the rest of this paper, we refer to models (and state-tracking mechanisms) as “AA-type” or “PAA-type” based on which cluster of signatures they exhibit. Results in Appendix C break down probe accuracies by sequence length, confirming that models solve sequences of exponentially longer length at deeper layers.
What exactly is the “non-parity residual” for models that learn PAA? We visualize the linear components of representations near the final layer(s) of PAA models trained on $S_{3}$. The representations of states can be cleanly decomposed into two orthogonal parts: the parity of the product and a residual cluster identity, forming a triangular prism. In Figure 4, we project representations from the PAA model for each of the six states onto these components. Even-parity states (darker colors) and odd-parity states (lighter colors) are symmetric. The three cluster “spokes” are spaced 60 degrees apart. [^3] Further details, including an analysis of $S_{5}$, can be found in Appendix D.
@@ 4.4 Generalization by Sequence Length
Figure 5: Generalization curves showing state and parity prediction accuracy as sequence lengths vary. We plot generalization curves for AA and PAA models on $S_{3}$ and $S_{5}$. In each plot, we show the 98% cutoff threshold, the sequence length at which accuracy dips below 98%. In the models that learned PAA, the parity cutoff is larger than the state cutoff, while in models that learned AA, the parity cutoff equals the state cutoff. Generally speaking, models that learned AA generalize better than ones that learned PAA.
We next evaluate how AA- and PAA-type models generalize to sequences of varying lengths. We evaluate whether LMs are able to generate states that are (1) exactly correct (“state accuracy”) and (2) have the correct parity (“parity accuracy”) for held-out inputs of varying length. In general, we find that models learn to generalize perfectly to sequences of up to a particular length (specifically, close to the length of their training data), then face a steep accuracy dropoff after a particular length (which we refer to as the “cutoff length”), rather than generalizing uniformly across all sequence lengths.
In Figure 5, where we plot the cutoff lengths at which each accuracy dips below 98%. We find that for models that learn PAA, the parity accuracy cutoff length is longer than the state accuracy cutoff length, whereas, for models that learn an AA-type mechanism, the parity accuracy cutoff length is equal to the state accuracy cutoff length. Furthermore, models that learn an AA-type mechanism tend to generalize better overall.
@@ 4.5 Attention Patterns
We look at attention patterns of LMs and check whether they can be used to differentiate between PAA models and AA models. Specifically, we find that in the early layers, PAA models exhibit parity heads, heads that place attention to odd-parity actions. Recall that the parity of state can be determined by counting the number of odd-parity actions, and taking the parity of the count (Equation 1). Examples of the parity head attention pattern are shown in Figure 13. We find no evidence of parity heads in any layer of AA models. See Appendix E for a formal metric measuring how much an attention head behaves like a parity head. We also find evidence that attention patterns in AA models sparsify in later layers of the network, forming a tree-like pattern expected of AA shown in Figure 14.
@@ 5.1 When in Training Do Distinct Mechanisms Arise?
We find that an LM’s eventual mechanism can be identified very early in training, based on the pattern of prediction errors. LMs that eventually learn AA improve the quality of their parity and state predictions in lockstep, while LMs that learn PAA learn in two phases: first they converge on learning the parities of states over the entire length of the training sequence; and only then learn to accurately predict the state itself. See Figure 6. Experimental details can be found in Appendix F.
Figure 6: Annotated training curves for models that learn the AA and PAA algorithms. In PAA models (blue), we find that convergence happens in two phases: in the first phase, they learn to generalize parities up to sequence length 100, and in the second, they learn to generalize the states up to sequence length. In AA models (orange), parities and states are learned simultaneously. Note that AA models also tend to converge faster to (ultimately) a lower loss than PAA models.
@@ 5.2 What Factors Affect Which Mechanism is Learned?
Whether an LM learns AA or PAA is a deterministic function of four factors: model architecture, size, initialization scheme, and training data. Our next experiments evaluate each of these factors in turn. We explore two different model architecture families of various sizes (GPT-2, Radford et al., 2019, and Pythia, Biderman et al., 2023), several different model initializations (pretrained and trained from scratch with different random initializations), and up to 12 different data ordering seeds.
We find that model architecture and initialization, rather than model size, are the biggest determining factors of what mechanism the model chooses to learn. Figure 7 shows the ratio of LMs that learn each mechanism, aggregated by model architecture and initialization. The low variance indicates a minimal effect of model size. [^4] A finer-grained breakdown of the ratio over each model size can be found in Figure 15. GPT-2 models, regardless of whether they are pretrained, are split roughly evenly between the two mechanisms, while Pythia models tend to learn AA when pre-trained and PAA when not.
@@
5.3 How Does Pretraining Affect Mechanism Choice?
@@
[…]
As shown in Figure 7, when from-scratch LMs are trained with our topic modeling NTP objective, they always learn an AA-type mechanism.
@@
We first train the entire model on predicting the state parity of the sequence to output token 1 if odd and 0 if even, before transitioning to training on the actual $S_{3}$ objective. In Figure 7, we show that we can induce GPT-2 and Pythia models to learn PAA when trained from scratch on parity. Notably, from-scratch Pythia models already exhibit a tendency to learn PAA as a baseline behavior. Therefore, we also apply this curriculum on pretrained Pythia models and find that it consistently converts the mechanism learned from AA-type to PAA-type. [^5] We discuss another parity curriculum using an extra parity loss term in Section G.3.
@@
As a control, we train LMs to predict a random next token in $S_{3}$, by generating length-100 documents of random $S_{3}$ elements sampled from a uniform distribution. We confirm that the control fine-tuning did not affect the ratio with which LMs learned each mechanism.
@@ 6 Conclusion
We have shown that LMs trained trained or fine-tuned on permutation tracking tasks tasks learn one of two distinct mechanisms: one consistent with an “associative algorithm” (AA) that composes action subsequences in parallel across successive layers; and one consistent with a “parity-associative algorithm” (PAA) which first computes a shallow parity heuristic in early layers, then computes a residual to the parity using an associative procedure. LMs that learn an AA-type mechanism tend to generalize better and converge faster; different choices of model architecture and pre-training scheme encourage the discovery of one mechanism or the other.
While a large number of other state tracking tasks can be reduced to the more complex permutation task we study ($S_{5}$), our experiments leave open the question of whether the specific mechanisms LMs use to solve $S_{5}$ are also deployed for these other tasks—including, most importantly, for modeling situations described in natural language.
@@ Acknowledgments
This work was supported by the OpenPhilanthropy foundation, the MIT Quest for Intelligence, and the National Science Foundation under grant IIS-2238240. BZL is additionally supported by a Clare Boothe Luce fellowship, and JDA is supported by a Sloan fellowship. This work benefited from many conversations during the Simons Institute Program on Language Models and Transformers. The authors would also like to thank Reuben Stern, Sebastian Zhu, and Gabe Grand for feedback on drafts of the paper.
@@ Appendix A A Constant-Depth Algorithm Exists for $S_{3}$
$S_{5}$ is the smallest non-solvable permutation group. $S_{3}$ is isomorphic to $D_{3}$, the symmetry group of an equilateral triangle, which can be generated by a transposition $a=(23)$ and a 3-cycle $b=(123)$. [^6] https://proofwiki.org/wiki/Symmetric_Group_on_3_Letters_is_Isomorphic_to_Dihedral_Group_D3 These generators satisfy the relation $ab=ba^{-1}$, which allows any word problem in $S_{3}$ to be reduced by tracking (1) the cumulative parity of transpositions and (2) the count of 3-cycles modulo 3. Since both parity checking and modular counting can be computed using constant-depth threshold circuits, the word problem for $S_{3}$ belongs to $\mathsf{TC}^{0}$.
@@ Appendix B Full Activation Patching Results
    In suffix patching, all tokens starting from $t$ up to one before the last token of the sequence ($h_{t:|A|-1,l}$) are patched at a particular layer $l$ [^7] We do not patch the last token as the last token residual contains the current accumulated information necessary for computing the final product, and almost always will destroy the prediction when patched, making this method uninformative as localization tool. We wish to see what other tokens the final token uses when constructing its own representation.
@@
Figure 8: Activation patching results across different types of localization (prefix, suffix, window patching) and different types of patching content (substitution, deletion). (A): Prefix-substitution results on only sequences with the same parity. We see the same exponential patching pattern in both AA and PAA models, showing that parity complements are computed in the same associative manner. (B): Prefix-substitution results on only sequences with opposite parities. We see the exponential patching pattern in AA models, meaning in AA models, parities are computed with the state. In PAA models, however, the patching pattern is roughly parallel, meaning parities are computed roughly in parallel in early layers. (C): Suffix-deletion results show that we can ignore progressively longer sequences of suffixes as we go down the layers of the network, consistent with how we believe AA and PAA work. (D): Window-deletion results show that important activations are arranged hierarchically, again consistent with how we believe AA and PAA work.
@@
Results are shown in Figure 8A (for same parity) and 8B (for opposite parity). We find that in AA models, parities are computed with the state – with both the same-parity and opposite-parity patching patterns displaying the same exponential curve. However, in PAA models, the patching patterns for same and opposite parities differ drastically. When parities are the same, only the parity complement must be computed to infer the final state; the patching pattern in this case indicates that the parity complement is computed in an associative manner. When the parities are different, the patching signature has a component that resembles a parallel patching signature, which is where the parity is computed. Restoring prefixes of layers before the parity is computed results in the entire prediction being of the correct parity, while restoring prefixes of layers after that results in the entire prediction being of the incorrect parity. We see that parities are computed roughly in parallel at early layers (around layers 3-5). Note that there is a middle region where restoring the prefixes shifts the logits towards the correct prediction, but not 100%: when prefixes in these regions are restored, the LM does not know the parity of the final answer, but does know some aspects of the parity complement, which was computed in an associative manner.
@@
Important activations are arranged in a hierarchically

### v3

@@ 4 What Mechanisms do Transformers Learn?
Figure 2: Activation patching on the residual stream for various Pythia models trained on $S_{3}$ and $S_{5}$. Each cell at layer $l$ and token $t$ represents the probability of the correct final state when the entire prefix up to $t$ at layer $l$ is restored. We find signatures matching the AA and PAA algorithms from Figure 1, with both models ignoring exponentially longer prefixes as we traverse down the layers, and PAA models containing intermediate representations that encode some information about the final state, but not its parity.
Nevertheless, our experiments successfully rule out some possible state tracking mechanisms and identify algorithmic features likely to be shared between the idealized mechanisms above and the true behavior learned by transformers. Specifically, our experiments yield evidence consistent with the associative algorithm (AA) in some models and the parity-associative algorithm (PAA) in other models, across architectures, sizes, and initializations.
@@ 4.1 Experimental Setup
We generate 1 million unique length-100 sequences of permutations in both $S_{3}$ and $S_{5}$. We split the data 90/10 for training/analysis, and fine-tune these models (using a cross-entropy loss) to predict the state corresponding to each prefix of each action sequence:
@@
where $p_{\texttt{LM}}\left(s_{n}\mid a_{0}\ldots a_{t}\right)$ is the probability the language model places on state token $s_{n}$ when conditioned on the length-$n$ prefix of the document.
@@
Figure 3: Accuracy of state probe and state parity probe across layers on $S_{3}$ and $S_{5}$ models sometimes match signatures for AA, and sometimes PAA. In all models, the state probe accuracy increases roughly exponentially with model depth. We find that in PAA models, the parity of the state is linearly decodable from earlier intermediate layers, while in the AA models shown above, the parity is never linearly encoded in any layer of the model. (In other AA models, the parity can only be linearly decoded at the final layer.)
Figure 4: In models that learn PAA on $S_{3}$, representations of the final product can be geometrically decomposed into two orthogonal directions, corresponding to the parity of the product (represented as the Z-axis in the above graph) and cluster identity of the product (represented by the X-Y plane). Note that the clusters are at 60 degrees to each other, and products of different parities within a cluster are equidistant from each other, with odd-parity products in one plane, and even-parity products in another plane.
@@ 4.2 Activation Patching
For both the $S_{3}$ and $S_{5}$ tasks, across training runs, we find that activation patching results exhibit two broad clusters of behavior. For some trained models, they match the activation patching signature associated with AA; in others, they match the signature of PAA—even when the only source of variability across training runs is the order in which data is presented. Results for prototypical AA- and PAA-type models, on both $S_{3}$ and $S_{5}$, are shown in Figure 2. Additional patching results in Appendix B confirm that patching intermediate representations of PAA-type models (the light-colored cells in Figure 2) results specifically in predictions with incorrect parity. We find standard deviations to be low in Figure 9, confirming the robustness of our signatures.
@@ 4.3 Probing
Test set accuracies of linear probes across LM layers $l$ are plotted in Figure 3. We report standard deviations of these accuracies in Table 1, which are all less than $10^{-3}$. We again find empirical signatures consistent with those predicted by AA and PAA, on both $S_{3}$ and $S_{5}$. Models with AA-type probing signatures always have AA-type patching signatures, and vice versa. Throughout the rest of this paper, we refer to models (and state-tracking mechanisms) as “AA-type” or “PAA-type” based on which cluster of signatures they exhibit. Results in Appendix C break down probe accuracies by sequence length, confirming that models solve sequences of exponentially longer length at deeper layers.
What exactly is the “non-parity residual” for PAA models? We visualize the linear components of representations near the final layer(s) of PAA models trained on $S_{3}$. The representations of states can be cleanly decomposed into two orthogonal parts: the parity of the product and a residual cluster identity, forming a triangular prism. In Figure 4, we project representations from the PAA model for each of the six states onto these components. Even-parity states (darker colors) and odd-parity states (lighter colors) are symmetric. The three cluster “spokes” are spaced 60 degrees apart. [^2] Further details, including an analysis of $S_{5}$, can be found in Appendix D.
@@ 4.4 Generalization by Sequence Length
Figure 5: Generalization curves showing state and parity prediction accuracy as sequence lengths vary. Models are trained on length-100 sequences and asked to generalize to varying lengths of sequences. We plot generalization curves for AA and PAA models on $S_{3}$ and $S_{5}$. In each plot, we show the 98% cutoff threshold, the sequence length at which accuracy dips below 98%. In the models that learned PAA, the parity cutoff is larger than the state cutoff, while in models that learned AA, the parity cutoff equals the state cutoff. Generally speaking, models that learned AA generalize better than ones that learned PAA.
We next evaluate the state and parity accuracy of AA- and PAA-type models for held-out inputs of varying length. In general, we find that models learn to generalize perfectly to sequences of up to the length of their training data, then face a steep accuracy dropoff after (which we refer to as the “cutoff length”), rather than generalizing uniformly across all sequence lengths.
In Figure 5, where we plot the cutoff lengths at which each accuracy dips below 98%. We find that for models that learn PAA, the parity accuracy cutoff length is much longer than the state accuracy cutoff length, whereas, for models that learn an AA-type mechanism, the two cutoff lengths are equal. Furthermore, models that learn an AA-type mechanism tend to generalize better overall.
Figure 6: Annotated training curves for models that learn the AA and PAA algorithms. In PAA models (blue), we find that convergence happens in two phases: in the first phase, they learn to generalize parities up to sequence length 100, and in the second, they learn to generalize the states. In AA models (orange), parities and states are learned simultaneously. Note that AA models also tend to converge faster to (ultimately) a lower loss than PAA models.
@@ 4.5 Attention Patterns
We look at attention patterns of LMs and check whether they can be used to differentiate between PAA models and AA models. Specifically, we find that in the early layers, PAA models exhibit parity heads, heads that place attention to odd-parity actions. Recall that the parity of a state can be determined by counting the number of odd-parity actions, and taking the parity of the count (Equation 1). Examples of the parity head attention pattern are shown in Figure 15. We find no evidence of parity heads in any layer of AA models. (See Appendix F for a formal metric measuring how much an attention head behaves like a parity head.) We also find evidence that attention patterns in AA models sparsify in later layers of the network, forming a tree-like pattern expected of AA, shown in Figure 16.
@@ 5.1 When in Training Do Distinct Mechanisms Arise?
We find that an LM’s eventual mechanism can be identified very early in training, based on the pattern of prediction errors. As in Figure 6, LMs that eventually learn AA improve the quality of their parity and state predictions in lockstep, while LMs that learn PAA learn in two phases: they first converge on learning parity over the entire length of the training sequence; and only then do they learn to accurately predict the state itself. [^3] Experimental details can be found in Appendix G.
@@ 5.2 What Factors Affect Which Mechanism is Learned?
Whether an LM learns AA or PAA is a deterministic function of four factors: model architecture, size, initialization scheme, and fine-tuning data order. Our next experiments evaluate each of these factors in turn. We explore two different model architecture families of various sizes (GPT-2, Radford et al., 2019, and Pythia, Biderman et al., 2023), several different model initializations (pre-trained on the Pile and trained from scratch with different random initializations), and up to 12 different data ordering seeds.
We find that model architecture and initialization, rather than model size, are the biggest determining factors of what mechanism the model chooses to learn. Figure 7 shows the ratio of LMs that learn each mechanism, aggregated by model architecture and initialization. The low variance indicates a minimal effect of model size. [^4] A finer-grained breakdown of the ratio over each model size can be found in Figure 17. GPT-2 models, pre-trained or not, are split roughly evenly between the two mechanisms, while Pythia models tend to learn AA when pre-trained on the Pile, and PAA when not.
@@
5.3 How Does Pre-training Affect Mechanism Choice?
@@
As a controlled way of studying how the next-token-prediction (NTP) objective affects which mechanism LMs converge to, we generate length-100 documents with only $S_{3}$ elements as vocabulary items, and pre-train (randomly initialized) LMs with NTP on these documents, before training them on $S_{3}$. Specifically, the documents are generated from a topic model with parameters: # $\text{\# of topics }=4$, $\alpha=0.3$, and $\beta=0.1$, where $\alpha$ is the density of topics in each document and $\beta$ is density of words in each topic. As shown in Figure 7, when from-scratch LMs are trained with our topic modeling NTP objective, they always learn an AA-type mechanism.
@@
We first train the entire model on predicting the state parity of the sequence to output token 1 if odd and 0 if even, before transitioning to training on the actual $S_{3}$ objective. In Figure 7, we show that we can induce GPT-2 and Pythia models to learn PAA when trained from scratch on parity. Notably, from-scratch Pythia models already tend to learn PAA as a baseline behavior. Therefore, we also apply this curriculum on Pythia models pre-trained on the Pile, and find that it consistently converts the mechanism learned from AA-type to PAA-type. [^5] We discuss another parity curriculum using an extra parity loss term in Section H.3.
@@
As a control, we train LMs on length-100 documents of random $S_{3}$ elements sampled from a uniform distribution. We confirm that the control fine-tuning did not affect the ratio with which LMs learned each mechanism.
@@ 6 Conclusion
We have shown that LMs trained on permutation tracking tasks learn one of two distinct mechanisms: one consistent with an “associative algorithm” (AA) that composes action subsequences in parallel across successive layers; and another with a “parity-associative algorithm” (PAA) which first computes a shallow parity heuristic in early layers and then computes a residual to the parity using an associative procedure. LMs that learn an AA-type mechanism tend to generalize better and converge faster; different choices of model architecture and training scheme encourage the discovery of one mechanism over another.
While a large number of other state tracking tasks can be reduced to the more complex permutation task we study ($S_{5}$), our experiments leave open the question of whether the specific mechanisms LMs use to solve $S_{5}$ are also deployed for these other tasks.
@@ Acknowledgments
This work was supported by the OpenPhilanthropy foundation, the MIT Quest for Intelligence, and the National Science Foundation under grant IIS-2238240. BZL is additionally supported by a Clare Boothe Luce fellowship, and JA is supported by a Sloan fellowship. This work benefited from many conversations during the Simons Institute Program on Language Models and Transformers. The authors would also like to thank Reuben Stern, Sebastian Zhu, and Gabe Grand for feedback on drafts of the paper.
@@ Appendix A A Constant-Depth Algorithm Exists for $S_{3}$
$S_{5}$ is the smallest non-solvable permutation group. $S_{3}$ is isomorphic to $D_{3}$, the symmetry group of an equilateral triangle, which can be generated by a transposition $a=(23)$ and a 3-cycle b=(123)b=(123). [^6] https://proofwiki.org/wiki/Symmetric_Group_on_3_Letters_is_Isomorphic_to_Dihedral_Group_D3 These generators satisfy the relation $ab=ba^{-1}$, which allows any word problem in $S_{3}$ to be reduced by tracking (1) the cumulative parity of transpositions and (2) the count of 3-cycles modulo 3. Since both parity checking and modular counting can be computed using constant-depth threshold circuits, the word problem for $S_{3}$ belongs to $\mathsf{TC}^{0}$.
@@ Appendix B Full Activation Patching Results
    In suffix patching, all tokens starting from $t$ up to one before the last token of the sequence ($h_{t:|A|-1,l}$) are patched at a particular layer ll. [^7] We do not patch the last token as the last token residual contains the current accumulated information necessary for computing the final product, and almost always will destroy the prediction when patched, making this method uninformative as localization tool. We wish to see what other tokens the final token uses when constructing its representation.
@@
Figure 8: Activation patching results across different types of localization (prefix, suffix, window patching) and different types of patching content (substitution, deletion). (A): Prefix-substitution results on only sequences with the same parity. We see the same exponential patching pattern in both AA and PAA models, showing that parity complements are computed in the same associative manner. (B): Prefix-substitution results on only sequences with opposite parities. We see the exponential patching pattern in AA models, meaning that in AA models, parities are computed with the state. In PAA models, however, the patching pattern is roughly parallel, meaning parities are computed roughly in parallel in early layers. (C): Suffix-deletion results show that we can ignore progressively longer sequences of suffixes as we go down the layers of the network, consistent with how we believe AA and PAA work. (D): Window-deletion results show that important activations are arranged hierarchically, again consistent with how we believe AA and PAA work.
@@
Results are shown in Figure 8A (for same parity) and 8B (for opposite parity). We find that in AA models, parities are computed with the state – with both the same-parity and opposite-parity patching patterns displaying the same exponential curve. However, in PAA models, the patching patterns for same- and opposite-parities differ drastically. When parities are the same, only the parity complement must be computed to infer the final state; the patching pattern in this case indicates that the parity complement is computed in an associative manner. When the parities are different, the patching signature has a component that resembles a parallel patching signature, which is where the parity is computed. Restoring prefixes of layers before the parity is computed results in the entire prediction being of the correct parity, while restoring prefixes of layers after that results in the entire prediction being of the incorrect parity. We see that parities are computed roughly in parallel at early layers (around layers 3-5). Note that there is a middle region where restoring the prefixes shifts the logits towards the correct prediction, but not 100%: when prefixes in these regions are restored, the LM does not know the parity of the final answer, but does know some aspects of the parity complement, which was computed in an associative manner.
@@
Important activations are arranged hierarchically
@@
Figure 9: Standard deviations of the prefix substitution activation patching results across 200 S3 input pairs, for (A) AA models, (B) PAA models on pairs with the same parity, and (C) PAA models on pairs with opposite parity. We find generally low standard deviations across examples.
Patching signatures are relatively consistent across examples
In Figure 9, we plot the standard deviations across 200 pairs of S3 inputs for the following three sets of results: (A) prefix substitution patching of AA models, (B) prefix substitution patching of PAA models (opposite parity inputs), (C) prefix substitution patching of PAA models (same parity inputs). We find relatively low standard deviations in all three cases, showing that these signatures hold across different examples.